A and B are 420 m apart. A RUNS A CONSTANT VELOCITY OF 9M/ S. A runs with a constant velocity of 5m/s.

Where and when do they meet?


1
Expert's answer
2021-01-19T07:12:11-0500

The relative velosity of A with respect to B is

"v_{AB}=v_A+v_B=\rm 9\: m/s+5\: m/s=14\: m/s"

The A and B will meet after

"\Delta t=\frac{d}{v_{AB}}=\rm\frac{420\: m}{14\: m/s}=30\: s"

The A will travel

"d_A=v_A\Delta t=\rm 9\: m/s\cdot 30\: s=270\: m"